The ages of Ernie and Bert are consecutive odd integers. In six years, the sum of their ages will be four times Ernie's present age. How old is Ernie if Bert is older than Ernie?
Answer by richwmiller(17219)

e+b+12=4e
b=e+2
2e+14=4e
e=7
b=9
check
7+6+9+6=4*7
13+15=28
28=28
ok